# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v6

@@ 5 SMD Maintenance Mechanisms
We propose SMD-based maintenance mechanisms for three use cases. However, SMD is not limited to these three use cases and it can be used to support more operations in DRAM.
@@ 5.1 Use Case 1: DRAM Refresh
In conventional DRAM, the MC periodically issues REF commands to initiate a DRAM refresh operation. This approach is inefficient due to two main reasons. First, transmitting REFs over the DRAM command bus within the refresh period (e.g., 64, 32, or even depending on the refresh rate) consumes energy and increases the command bus utilization. Due to transmitting a REF over the command bus, the MC may delay a command to another rank in the same channel, which increases DRAM access latency [16]. Second, an entire bank becomes inaccessible while being refreshed although a REF command refreshes only a few rows in the bank. This incurs up to 50% loss in memory throughput [17].
Leveraging SMD, we design a new efficient maintenance mechanism for DRAM refresh: Fixed-Rate Refresh.
@@
Fixed-Rate Refresh (SMD-FR). SMD-FR refreshes DRAM rows in a fixed time interval (i.e., tREFI), similar to conventional DRAM refresh. To limit the time that the MC waits for a region to get unlocked, SMD-FR refreshes (Refresh Granularity) number of rows from a lock region and switches to refreshing another region.
@@
SMD-FR (Fig. 4) operates independently in each DRAM bank and it uses three counters for managing the refresh operations: pending refresh counter, lock region counter, and row address counter. The pending refresh counter is initially zero and SMD-FR increments it by one at the end of a tREFI interval . SMD-FR allows up to refresh operations to be accumulated in the pending refresh counter.666DDR4 [78, 83, 85] allows the MC to postpone issuing up to REF commands in order to serve pending memory requests first. Because the MC can keep a row open for a limited time (section 4.3), the pending refresh counter never exceeds the value .
@@
The lock region and row address counters indicate the next row to refresh. When pending refresh counter is greater than zero, SMD-FR attempts to lock the region indicated by the lock region counter every clock cycle until it successfully locks the region . In some cycles, SMD-FR may fail to lock the region either because the lock region contains an active row or the region is locked by another maintenance mechanism. When SMD-FR successfully locks the region, it initiates a refresh operation that refreshes number of rows in the lock region starting from the row indicated by the row address counter . The refresh operation is conducted by activating and precharging a row to restore the charge in its DRAM cells. We empirically find to be a favorable design point.
@@
After the refresh operation completes, SMD-FR releases the locked region and increments only the lock region counter. When the lock region counter rolls back to zero, SMD-FR also increments by one the row address counter . Finally, SMD-FR decrements the target refresh counter .
@@
Variable Refresh (SMD-VR). In conventional DRAM, all rows are uniformly refreshed with the same refresh period. However, the actual data retention times of different rows in the same DRAM chip greatly vary mainly due to manufacturing process variation and design-induced variation [17, 15, 86, 38, 1, 18]. In fact, only hundreds of “weak” rows across an entire DRAM chip require to be refreshed at the default rate, and a vast majority of the rows can correctly operate when the refresh period is doubled or quadrupled [17]. Eliminating unnecessary refreshes to DRAM rows that do not contain weak cells can significantly mitigate the performance and energy consumption overhead of DRAM refresh [17].
@@
The key idea of SMD-VR is to group DRAM rows into multiple retention time bins and refresh a row based on the bin that it belongs to. To achieve low design complexity, SMD-VR uses only two bins: 1) retention-weak rows that have retention time less than and 2) rows that have retention time more than . Inspired by RAIDR [17], SMD-VR stores the addresses of retention-weak rows using a per-bank Bloom Filter [87], which is a space-efficient probabilistic data structure for representing set membership. We assume retention-weak rows are already inserted into the Bloom Filters by the DRAM vendors during post-manufacturing tests.777Alternatively, SMD can be used to develop a maintenance mechanism that performs retention profiling. We leave the development and analysis of it to future work.
@@
Unless stated otherwise, we assume and . Therefore, SMD-VR refreshes the DRAM rows that are not in the Bloom Filter once every four consecutive refresh periods as these rows can retain their data correctly for at least four refresh periods.
After determining which DRAM rows need refresh, SMD-VR operates in a way similar to SMD-FR.
@@ 5.2 Use Case 2: RowHammer Protection
Repeatedly activating and precharging (i.e., hammering) a (aggressor) DRAM row causes RowHammer errors in the cells of a nearby (victim) DRAM row [88, 89, 90, 91, 92, 93, 94, 95]. DRAM manufacturers equip their existing DRAM chips with in-DRAM RowHammer protection mechanisms, generally referred to as Target Row Refresh (TRR) [74, 75, 76]. At a high level, TRR protects against RowHammer by detecting an aggressor row and refreshing its victim rows. Because a conventional DRAM chip cannot initiate a refresh operation by itself, TRR refreshes victim rows by taking advantage of the slack time available in the refresh latency (i.e., tRFC), originally used to perform only periodic DRAM refresh [74, 75, 76]. Recent works [74, 75, 76, 96] demonstrate a variety of new RowHammer access patterns that circumvent the TRR protection in chips of all three major DRAM vendors, proving that the existing DRAM interface is not well suited to enable strong RowHammer protection especially as RowHammer becomes a bigger problem with DRAM technology scaling. We use SMD to develop a new maintenance mechanism that overcomes the limitations of the existing TRR mechanisms by initiating victim row refresh within the DRAM chip.
Probabilistic RowHammer Protection (PRP). Inspired by PARA [10], we implement an in-DRAM maintenance mechanism called Probabilistic RowHammer Protection (SMD-PRP). The high level idea is to refresh the nearby rows of an activated row with a small probability. PARA is proposed as a mechanisms in the MC, which makes it difficult to adopt since victim rows are not always known to the MC. SMD enables us to overcome this issue by implementing the PARA-inspired SMD-PRP mechanism completely within the DRAM chip. In addition, SMD-PRP avoids explicit ACT and PRE commands to be sent over the DRAM bus. We omit the implementation details of SMD-PRP due to space limitations.
Fig. 6 illustrates the operation of SMD-PRP.
@@
SMD-PRP with Aggressor Row Detection (SMD-PRP+). SMD-PRP refreshes victim rows with a small probability on every row activation, even does so for a row that has been activated only a few times. This results in unnecessary victim refreshes, especially for high values that strengthen the RowHammer protection as DRAM cells become more vulnerable to RowHammer with technology scaling.
We propose SMD-PRP+, which detects potential aggressor rows and probabilistically refreshes only their victim rows. The key idea of SMD-PRP+ is to track frequently-activated rows in a DRAM bank and refresh the neighbor rows of these rows using the region locking mechanism of SMD.
SMD-PRP+ tracks frequently-activated rows, within a rolling time window of length , using two Counting Bloom Filters (CBF) [97] that operate in time-interleaved manner. CBF is a Bloom Filter variant that represents the upperbound for the number of times an element is inserted into the CBF. We refer the reader to prior work for background on CBFs [54, 98].
Depicted in Fig. 7, SMD-PRP+ operation is based on 1) detecting a row that has been activated more than times within the most recent interval and 2) refreshing the neighbor rows of this row. must be set according to the minimum hammer count needed to cause a RowHammer bit flip in a given DRAM chip. A recent work shows that activations can cause bit flips in LPDDR4 chips [9]. We conservatively set to . must be equal to or larger than the refresh period in order to track all activations that happen until rows get refreshed by regular refresh operations. We set equal to the refresh period.
Initially, one of the CBFs is in active mode while the other is in passive mode. When activating a row, SMD-PRP+ inserts the address of the activated row to both CBFs . Then, SMD-PRP+ tests the active CBF to check if the accumulated insertion count exceeds the threshold . If so, SMD-PRP+ marks the row in the Marked Rows Table to refresh its neighbor rows with the probability of . SMD-PRP+ does not always mark the row because it is impossible to reset only the counters that correspond to the marked row address in the CBFs. If always marked, a subsequent activation of the same row would again cause the row to be marked, leading to unnecessary neighbor row refresh until all CBF counters are reset, which happens at the end of the interval. After a row is marked, steps - are same as steps - in SMD-PRP. Finally, at the end of an interval, SMD-PRP+ clears the active CBF and swaps the two CBFs to continuously track row activations within the most recent window .
Deterministic RowHammer Protection (DRP). We use SMD to implement SMD-DRP, a deterministic RowHammer protection based on the Graphene [48] mechanism keeps track of frequently activated DRAM rows. Different from Graphene, we implement SMD-DRP completely within a DRAM chip, whereas Graphene requires the MC to issue neighbor row refresh operations.888 We refer the reader to [48] for more details on the Graphene mechanism and its security proof. The operation of SMD-DRP is similar to the operation of Graphene with the difference of SMD-DRP being implemented completely within a DRAM chip, which does not affect the underlying operation, and thus the security proof of Graphene applies to SMD-DRP.
SMD-PRP and SMD-PRP+ are probabilistic mechanisms that provide statistical security guarantees against RowHammer attacks with being the security parameter. On an extremely security-critical system, a deterministic RowHammer protection mechanisms that guarantees the mitigation of RowHammer bit flips at all times may be desired.
@@
When SMD-DRP receives an ACT, it checks if the activated row address () exists in CT . If so, SMD-DRP increments the corresponding CT counter by one . Otherwise, SMD-DRP finds the smallest counter value () in CT and compares it to the value of the spillover counter (SP) , which is initially zero. If SP is equal to , SMD-DRP replaces the row address corresponding to in CT with and increments the corresponding CT counter by one . If SP is smaller than , SMD-DRP increments SP by one . When a CT counter is incremented in step , SMD-DRP checks if the counter value is a multiple of , which is the maximum number of times a row can be activated without refreshing its neighbors. If so, SMD-DRP refreshes the neighbors of . To prevent the counters from overflowing, SMD-DRP resets the CT counters and SP on every tREFW interval.
@@
where is the maximum number of activations that the MC can perform within a tREFW interval in a single bank. In our evaluations, we set to show that the overheads of SMD-DRP are small even for DRAM chips that are extremely vulnerable against RowHammer.
@@ 5.3 Use Case 3: Memory Scrubbing
To mitigate the increasing bit errors mainly caused by the continued DRAM technology scaling, DRAM vendors equip their DRAM chips with on-die Error Correction Codes (ECC) [99, 100, 101, 8]. On-die ECC is designed to correct a single bit error assuming that a failure mechanism is unlikely to incur more than one bit error in a codeword [100]. However, even when the assumption always holds, a failure mechanism can gradually incur two or more bit errors over time. A widely-used technique for preventing the accumulation of an uncorrectable number of bit errors is memory scrubbing. Memory scrubbing describes the process of periodically scanning the memory for bit errors in order to correct them before more errors occur.
@@
We propose SMD-based Memory Scrubbing (SMD-MS), which is an in-DRAM maintenance mechanism that periodically performs scrubbing on DRAM chips with on-die ECC.999To preserve ECC transparency and enable system-level policies that depend on ECC scrubbing information (e.g., detecting and replacing faulty DRAM chips), SMD-MS exposes critical ECC scrubbing information (e.g., the address of the DRAM row with the most corrected errors) to the memory controller via DRAM mode-status registers, similar to how ECC Error Check and Scrub does in DDR5 [72]. Compared to conventional MC-based scrubbing, SMD-MS eliminates moving data to the MC by scrubbing within the DRAM chip, and thus reduces the performance and energy overheads of memory scrubbing.
@@
SMD-MS operation resembles the operation of SMD-FR. Similar to SMD-FR, SMD-MS maintains pending scrub counter, lock region counter, and row address counter. SMD-MS increments the pending scrub counter at fixed intervals of . When the pending scrub counter is greater than zero, SMD-MS attempts to lock the region indicated by the lock region counter. After locking the region, SMD-MS performs scrubbing operation on the row indicated by the row address counter. The scrubbing operation takes more time than a refresh operation as performing scrubbing on a row consists of three time consuming steps for each codeword in the row: 1) reading the codeword, 2) performing ECC decoding and checking for bit errors, and 3) encoding and writing back the new codeword into the row only when the decoded codeword contains a bit error. Refreshing a row takes , whereas scrubbing a row takes when no bit errors are detected.101010We assume ECC decoding/encoding hardware is fully pipelined. In case of a bit error, writing a corrected codeword incurs latency. Therefore, SMD-MS keeps a region locked for longer than SMD-FR. When the scrubbing operation is complete, SMD-MS releases the lock region and increments the lock region and row address counters, and decrements the pending scrub counter as in SMD-FR.
Rank-Level Memory Scrubbing. Server memory modules typically use rank-level ECC [102, 103, 104]. We discuss why rank-level ECC and in-DRAM ECC scrubbing should be done separately but in a combined way in server modules. During a conventional DRAM RD operation, a DRAM chip with in-DRAM ECC (e.g., a DDR5 chip [72]) performs error correction for the in-DRAM ECC codeword without 1) writing the corrected codeword back and 2) keeping a record of the correction in mode-status registers [72]. Thus, in a rank-level ECC scrubbing operation, the memory controller cannot distinguish between an in-DRAM-ECC-corrected codeword and a codeword with no errors.
@@
In a computing system that uses memory modules with in-DRAM ECC, relying only on the rank-level ECC scrubbing to prevent the accumulation of errors is costly. This is because the memory controller must write back every codeword that it reads from DRAM for every rank-level ECC scrubbing operation. Doing so could incur performance overheads due to the long data bus turn-around latency in DRAM (e.g., in DDR5 is approximately [105, 72]) and energy overheads due to increased data movement between the memory controller and the DRAM chips. Instead, using SMD, the computing system could use both rank-level ECC scrubbing and in-DRAM ECC scrubbing to 1) leverage the (typically) stronger error correction capability of rank-level ECC and 2) leverage the performance and energy efficiency of in-DRAM ECC while preventing the accumulation of errors.

### v7

@@ 5 SMD Maintenance Mechanisms
We propose SMD-based maintenance mechanisms for three use cases. SMD is not limited to these three use cases and it can be used to support more operations in DRAM (as described in section 5.4 and section 5.5).
@@ 5.1 Use Case 1: DRAM Refresh
In conventional DRAM, the MC periodically issues REF commands to initiate a DRAM refresh operation. This approach is inefficient due to two main reasons. First, transmitting thousands of (e.g., ) REFs over the DRAM command bus within the refresh period (e.g., 64, 32, or even depending on the refresh rate [119]) consumes energy and increases the command bus utilization. A REF command may delay a command to another DRAM component (e.g., bank, bank group, rank) in the same channel, which increases DRAM access latency [19]. Second, an entire bank becomes inaccessible while being refreshed although a REF command refreshes only a small fraction of rows in the bank.
Leveraging SMD, we design two new efficient maintenance mechanisms for periodic DRAM refresh: Fixed-Rate Refresh and Variable-Rate Refresh. The Fixed-Rate Refresh mechanism aims to refresh all rows in a DRAM chip uniformly at a fixed refresh period, similar to conventional DRAM refresh. The Variable Refresh mechanism inspired by RAIDR [20] refreshes different DRAM rows at different intervals based on their retention time characteristics.
@@
Fixed-Rate Refresh (SMD-FR). SMD-FR refreshes DRAM rows in a fixed time interval (i.e., tREFI), similar to conventional DRAM refresh. To limit the time that the MC waits for a region to get unlocked, SMD-FR refreshes (Refresh Granularity) number of rows from a lock region and then switches to refreshing another region.
@@
SMD-FR (Fig. 4) operates independently in each DRAM bank and it uses three counters for managing the refresh operations: pending refresh counter, lock region counter, and row address counter. The pending refresh counter is initially zero and SMD-FR increments it by one at the end of every tREFI interval . SMD-FR allows up to N refresh operations to be accumulated in the pending refresh counter. We set N=8.999DDR4 [141, 153, 158] allows the MC to postpone issuing up to REF commands in order to serve pending memory requests first [19, 36]. DDR5 reduces this number to [119]. Because the MC can keep a row open for a limited time (section 4.3), the pending refresh counter never exceeds the value N.
@@
The lock region and row address counters indicate the next row to refresh. When pending refresh counter is greater than zero, SMD-FR attempts to lock the region indicated by the lock region counter every clock cycle until it successfully locks the region . In some cycles, SMD-FR may fail to lock the region either because the lock region contains an active row or the region is locked by another maintenance mechanism. When SMD-FR successfully locks the region, it initiates a refresh operation that refreshes number of rows in the lock region starting from the row indicated by the row address counter . We empirically find to be a favorable design point.101010We evaluate RG = 1, 2, 4, 8, 16, 32, and 64. We find that RG = 8 and RG = 4 yield the two highest performance improvements among all tested RG values. Making larger increases the latency of a single refresh operation, which prolongs the time that a lock region remains unavailable for access. In contrast, a too small causes switching from one lock region to another very often, increasing interference with accesses.
@@
After the refresh operation completes, SMD-FR releases the locked region and increments only the lock region counter. When the lock region counter rolls back to zero, SMD-FR also increments the row address counter . Finally, SMD-FR decrements the pending refresh counter .
5.1.1 Variable Refresh (SMD-VR)
@@
In conventional DRAM, all rows are uniformly refreshed with the same refresh period. However, the actual data retention times of different rows in the same DRAM chip greatly vary mainly due to manufacturing process variation and design-induced variation [20, 17, 161, 41, 1, 21]. In fact, only hundreds of “weak” rows across an entire DRAM chip require to be refreshed at the default rate, and a vast majority of the rows can correctly operate when the refresh period is doubled or quadrupled [20]. Eliminating unnecessary refreshes to DRAM rows that do not contain weak cells can significantly mitigate the performance and energy consumption overhead of DRAM refresh [20].
@@
The key idea of SMD-VR is to group DRAM rows into multiple retention time bins and refresh a row based on the bin that it belongs to. To achieve low design complexity, SMD-VR uses only two bins: 1) retention-weak rows that have retention time less than and 2) rows that have retention time more than . Inspired by RAIDR [20], SMD-VR stores the addresses of retention-weak rows using a per-bank Bloom Filter [162], which is a space-efficient probabilistic data structure for representing set membership. We assume retention-weak rows are already inserted into the Bloom Filters by the DRAM vendors during post-manufacturing tests.111111Alternatively, SMD can be used to develop a maintenance mechanism that performs retention profiling. We leave the development and analysis of it to future work.
@@
Unless stated otherwise, we assume and . Therefore, SMD-VR refreshes the DRAM rows that are not in the Bloom Filter once every four consecutive refresh periods as these rows can retain their data correctly for at least four refresh periods. After determining which DRAM rows need refresh, SMD-VR operates in a way similar to SMD-FR.
@@ 5.2 Use Case 2: RowHammer Protection
Repeatedly activating and precharging (i.e., hammering) a DRAM row (aggressor) causes RowHammer bitflips [12] in the cells of a nearby DRAM row (victim) [12, 15, 18, 163, 164, 165, 166, 167, 168, 169, 170]. DRAM manufacturers equip their existing DRAM chips with in-DRAM RowHammer protection mechanisms, generally referred to as Target Row Refresh (TRR) [66, 64]. At a high level, TRR protects against RowHammer by detecting an aggressor row and refreshing its victim rows. Because a conventional DRAM chip cannot initiate a refresh operation by itself, TRR refreshes victim rows by taking advantage of the slack time available in the refresh latency (i.e., tRFC), originally used to perform only periodic DRAM refresh [66, 64, 138]. Recent works (e.g., [66, 64, 138, 171]) demonstrate a variety of new RowHammer access patterns that circumvent the TRR protection in chips of all three major DRAM vendors, demonstrating that the existing DRAM interface is not well suited to enable strong RowHammer protection especially as RowHammer becomes a bigger problem with DRAM technology scaling [11, 64, 172, 173, 174, 175, 15, 18]. We adapt two existing RowHammer protection mechanisms (PARA [12] and Graphene [61]) to be implemented freely in DRAM. These adaptations overcome the limitations of the existing TRR mechanisms by initiating victim row refresh within the DRAM chip.
Probabilistic RowHammer Protection (PRP). Inspired by PARA [12], we implement an in-DRAM maintenance mechanism called Probabilistic RowHammer Protection (SMD-PRP). The high level idea is to refresh the nearby rows of an activated row with a small probability. PARA is proposed as a mechanisms in the MC, which makes it difficult to adopt since victim rows are not always known to the MC. SMD enables us to overcome this issue by implementing the PARA-inspired SMD-PRP mechanism completely within the DRAM chip. In addition, SMD-PRP avoids explicit ACT and PRE commands to be sent over the DRAM bus. Fig. 6 illustrates the operation of SMD-PRP.
@@
Deterministic RowHammer Protection. We use SMD to implement SMD-DRP, a deterministic RowHammer protection technique based on the Graphene mechanism [61] that keeps track of frequently activated DRAM rows. Different from Graphene, we implement SMD-DRP completely within a DRAM chip, whereas Graphene requires the MC to issue neighbor row refresh operations.121212 We refer the reader to [61] for more details on the Graphene mechanism and its security proof. The operation of SMD-DRP is similar to the operation of Graphene with the difference that SMD-DRP is implemented completely within a DRAM chip, which does not affect the underlying operation. Thus, the security proof of Graphene applies to SMD-DRP.
@@
When SMD-DRP receives an ACT, it checks if the activated row address () exists in CT . If so, SMD-DRP increments the corresponding CT counter by one . Otherwise, SMD-DRP finds the smallest counter value () in CT and compares it to the value of the spillover counter (SP) , which is initially zero. The key idea of SP is to track the activation count of all DRAM rows that are not actively tracked by any of the CT entries. SP’s value is an upper bound for the maximum number of activations any (not-actively-tracked) DRAM row received. If SP is equal to , SMD-DRP replaces the row address corresponding to in CT with and increments the corresponding CT counter by one . If SP is smaller than , SMD-DRP increments SP by one . When a CT counter is incremented in step , SMD-DRP checks if the counter value is a multiple of , which is the maximum number of times a row can be activated without refreshing its neighbors (determined based on the RowHammer threshold [61]). If so, SMD-DRP refreshes the neighbor rows of . To prevent the counters from overflowing, SMD-DRP resets the CT counters and SP on every tREFW interval.
@@
where is the maximum number of activations that the MC can perform within a tREFW interval in a single bank. In our evaluations, we set to show that the overheads of SMD-DRP are small even for DRAM chips that are significantly vulnerable to RowHammer.131313When the maintenance operation time (MOT) mode (section 4.5) is enabled, SMD-DRP is configured with a smaller to provide the same security guarantee as when the MOT mode is not enabled. The reduction in is calculated as the number of ACT commands that the memory controller can issue to an aggressor row outside MOT. For example, if the DRAM chip switches into MOT mode every 3.9 µs (i.e., the MOT_period is 3.9µs), should be 3900/45 (MOT_period/tRC) smaller than when the MOT mode is not enabled.
@@ 5.3 Use Case 3: Memory Scrubbing
To mitigate the increasing bit errors mainly caused by the continued DRAM technology scaling, DRAM vendors equip their DRAM chips with on-die Error Correction Codes (ECC) [176, 177, 178, 10, 179, 180, 181, 182, 183]. On-die ECC is designed to correct a single bit error assuming that a failure mechanism is unlikely to incur more than one bit error in a codeword [177]. However, even if the assumption always holds, a failure mechanism can gradually incur two or more bit errors over time. A widely-used technique for preventing the accumulation of an uncorrectable number of bit errors is memory scrubbing. Memory scrubbing describes the process of periodically scanning the memory for bit errors in order to correct them before more errors occur.
@@
We propose SMD-based Memory Scrubbing (SMD-MS), which is an in-DRAM maintenance mechanism that periodically performs scrubbing on DRAM chips with on-die ECC.141414To preserve ECC transparency and enable system-level policies that depend on ECC scrubbing information (e.g., detecting and replacing faulty DRAM chips), SMD-MS exposes critical ECC scrubbing information (e.g., the address of the DRAM row with the most corrected errors) to the memory controller via DRAM mode-status registers, similar to what ECC Error Check and Scrub does in DDR5 [119]. Compared to conventional MC-based scrubbing, SMD-MS avoids moving data to the MC by performing scrubbing completely within the DRAM chip, thereby reducing the performance and energy overheads of MC-based memory scrubbing.
@@
SMD-MS operation resembles the operation of SMD-FR. Similar to SMD-FR, SMD-MS maintains pending scrub counter, lock region counter, and row address counter. SMD-MS increments the pending scrub counter at fixed intervals of . When the pending scrub counter is greater than zero, SMD-MS attempts to lock the region indicated by the lock region counter. After locking the region, SMD-MS performs a scrubbing operation on the row indicated by the row address counter. The scrubbing operation takes more time than a refresh operation as performing scrubbing on a row consists of three time consuming steps for each codeword in the row: 1) reading the codeword, 2) performing ECC decoding and checking for bit errors, and 3) encoding and writing back the new codeword into the row only when the decoded codeword contains a bit error. Refreshing a row takes , whereas scrubbing a row takes when no bit errors are detected.151515We assume ECC decoding/encoding hardware is fully pipelined. In case of a bit error, writing a corrected codeword incurs latency. Therefore, SMD-MS keeps a region locked for much longer than SMD-FR. When the scrubbing operation is complete, SMD-MS releases the lock region and increments the lock region and row address counters, and decrements the pending scrub counter.
@@
Rank-Level Memory Scrubbing. Server memory modules typically use rank-level ECC [179, 184, 185, 130, 184, 178]. We discuss why rank-level ECC and in-DRAM ECC scrubbing should be done separately but in a combined way in server modules. During a conventional DRAM RD operation, a DRAM chip with in-DRAM ECC (e.g., a DDR5 chip [119]) performs error correction for the in-DRAM ECC codeword without 1) writing the corrected codeword back and 2) keeping a record of the correction in mode-status registers [119]. Thus, in a rank-level ECC scrubbing operation, the memory controller cannot distinguish between an in-DRAM-ECC-corrected codeword and a codeword with no errors. Therefore, the memory controller writes back every codeword that it reads from DRAM for every rank-level ECC scrubbing operation. Doing so could incur performance overheads due to the long data bus turn-around latency in DRAM (e.g., in DDR5 is approximately [186, 119]) and energy overheads due to increased data movement between the MC and the DRAM chips. Instead, using SMD, the computing system could use both rank-level ECC scrubbing and in-DRAM ECC scrubbing to 1) leverage the (typically) stronger error correction capability of rank-level ECC and 2) leverage the performance and energy efficiency of in-DRAM ECC while preventing the accumulation of errors.